# ATRO: A Fast Algorithm for Topology Engineering of Reconfigurable Datacenter Networks

### v2

@@ Abstract
Index Terms:
Reconfigurable Data Center Networks, Optical Circuit Switches, Topology Engineering, Traffic Engineering.
@@ I Introduction
Reconfigurable data center networks (DCNs) allow the logical topology to adapt to evolving traffic patterns, and two dominant architectural paradigms have emerged: multi-hop and one-hop configurations. Multi-hop designs, such as Jupiter Evolving [4], maintain a relatively static topology and rely on dynamic traffic engineering to route traffic over multiple paths. In contrast, one-hop designs reconfigure OCSs at each scheduling interval to directly connect communicating pods [5, 6]. To minimize scheduling latency, these systems often eliminate routing decisions entirely and assume direct paths [3]. As a result, topology optimization (TO) is the central challenge in one-hop settings, while joint topology and routing optimization (TRO) is essential in multi-hop networks.
@@
We extensively evaluate ATRO in both scenarios and find that it consistently outperforms previous methods in maximum link utilization (MLU) and runtime, achieving up to $10\times$ speed-ups while maintaining high solution quality. Its modular structure, scalability, and practical efficiency make it suitable for a wide range of deployment scenarios. To facilitate further research, we will release our code on GitHub.
@@ II-A Problem Formulation
Similarly to [7, 8], we consider a reconfigurable DCN with $N$ pods, modeled as a directed graph $G=(V,E)$, where $V$ is the set of pods and $E$ comprises all potential logical links. Each pod $i$ has $R_{i}$ ports, each of capacity $S_{i}$, and the capacity of any logical link $(i,j)$ is $S_{i,j}=\min(S_{i},S_{j})$. The traffic demand from pod $i$ to pod $j$ is denoted $D_{i,j}$. We assume traffic is routed via at most two hops, a constraint widely adopted in production-scale DCNs [4]. We define a path as a triad $(s,k,d)$, where $s$ is the source, $d$ is the destination, and $k$ is either an intermediate relay (if $k\neq d$) or the destination itself (if directly routed). The set of all permissible paths is denoted $\mathcal{P}$, and for each source-destination pair $(i,j)$, we define the set of valid relay options as $\mathcal{K}_{ij}=\{k\mid(i,k,j)\in\mathcal{P}\}$. We introduce $f_{i,j,k}$ to denote the fraction of $i\to j$ traffic routed through node $k$ (with $k=j$ representing direct routing). Our goal is to jointly optimize the integer-valued logical topology $n_{i,j}$} and the continuous routing variables $f_{i,j,k}$} to minimize the maximum link utilization $u$. The joint topology and routing optimization (TRO) problem is formulated as Equation 1.
Constraint (1b) limits the total number of logical links per pod by available port count. Constraint (1c) enforces link symmetry, and requires that $n_{i,j}\in\mathbb{Z}^{+}$, i.e., each logical link count must be a non-negative integer. Constraints (1d)–(1e) restrict routing to valid two-hop paths. Constraint (1f) bounds link load by capacity. While our framework naturally generalizes to longer path-based routing [8], we adopt the two-hop model for clarity and tractability.
@@ II-B Limitations of Relaxation-Based Approaches
Figure 1: Illustration of infeasibility in the rounded TRO solution produced by COUDER. Subfigure (b) shows a relaxed LP solution with fractional link allocations. Subfigure (c) demonstrates an infeasible rounding outcome where PoD A and B are disconnected. Subfigure (d) shows a feasible and optimal integer solution.
@@ II-C Key Insight and Theoretical Discussion
  • Modular decomposition. ATRO decouples topology and routing, allowing each subproblem to be solved independently using specialized techniques. For example, the TO component can be solved via our custom binary search algorithm (ABSM), while the RO component supports LP-based or solver-free traffic engineering methods.
@@
  • Estimable and typically small optimality gap. ATRO directly optimizes the original MINLP, allowing LP-relaxed solutions to serve as lower bounds. In practice, it often converges near-optimal solutions by letting TO reshape global topology and RO mitigate local congestion. This synergy yields strong performance. In rare cases involving sparse demands, ATRO retains some links that an exact solver would remove. This occurs not due to actual traffic requirements, but because the routing step lacks incentives to suppress load on any specific link to zero, preventing the TO step from pruning it. Although slightly suboptimal, this behavior improves robustness.
@@ IV-B One-Hop Evaluation (TO Component)
Stress Test. ATRO outperforms MILP-TO in one-hop TO problems, highlighting the need for speed in real-time topology reconfiguration despite MILP-TO’s capability to handle moderate sizes quickly. Stress testing with larger topologies (see Table II) shows ATRO maintains optimal MLU (see Fig.6(a)) and sub-second runtimes (see Fig.6(b)), unlike MILP-TO, which struggles with large-scale topology. These findings demonstrate ATRO’s scalability and suitability for latency-sensitive scenarios in one hop scenario.
@@
Link Count Analysis. In addition to MLU and runtime, we evaluate the number of logical links each method provisions to satisfy traffic demand. As seen in Fig. 7, ATRO consistently provisions the fewest links, thanks to theorem 1. MILP-TO often over-provisions due to solver rounding effects, leading to inefficiencies. MCF and BvN lack explicit link minimization objectives and hence result in inflated configurations, particularly in larger topologies like Topo 64 and 128.
@@ VI Conclusion
We present ATRO, a modular framework for computing logical topologies in reconfigurable data center networks (DCNs). In the general multi-hop setting, ATRO alternates between topology optimization (TO) and routing optimization (RO) using lightweight, scalable subroutines. The TO step is solver-free and solved optimally via our proposed Accelerated Binary Search Method (ABSM), while the RO step supports both LP solvers and TE accelerators—enabling fully solver-free execution if desired. Extensive experiments show that ATRO matches or outperforms existing baselines in both performance and runtime, achieving low-latency scheduling in one-hop settings and efficient scalability in large-scale, multi-hop scenarios. Its convergence-guaranteed, plug-and-play design with warm-start and hybrid support makes ATRO well-suited for real-time, dynamic DCNs.

### v3

@@ Abstract
Yingming Mao and Qiaozhu Zhai contributed equally to this work.
Corresponding author: Qiaozhu Zhai (qzzhai@sei.xjtu.edu.cn).
This work was supported by the Fundamental Research Funds for the Central Universities (xzy022025039) and by the Fundamental Research Funds for the Central Universities (xtr052025030).
@@ I Introduction
Reconfigurable DCNs adapt logical topologies to traffic patterns via two paradigms: multi-hop and one-hop. Multi-hop designs, e.g., Jupiter Evolving [4], rely on dynamic routing over relatively static structures. Conversely, one-hop designs [5, 6] reconfigure OCSs to directly connect Points of Delivery (PoDs), often bypassing routing decisions to minimize latency [3]. Consequently, topology optimization (TO) dominates one-hop settings, while joint topology and routing optimization (TRO) is vital for multi-hop networks.
@@
We extensively evaluate ATRO in both scenarios and find that it consistently outperforms previous methods in maximum link utilization (MLU) and runtime, achieving up to $10\times$ speed-ups while maintaining high solution quality. Its modular structure, scalability, and practical efficiency make it suitable for a wide range of deployment scenarios. To facilitate further research, our code is available at [19].
@@ II-A Problem Formulation
Similarly to [7, 8], we consider a reconfigurable DCN with $N$ P, modeled as a directed graph $G=(V,E)$, where $V$ is the set of Po and $E$ comprises all potential logical links. Each PoD $i$ has $R_{i}$ ports, each of capacity $S_{i}$, and the capacity of any logical link $(i,j)$ is $S_{i,j}=\min(S_{i},S_{j})$. The traffic demand from PoD $i$ to PoD $j$ is denoted $D_{i,j}$. Let $n_{i,j}\in\mathbb{Z}_{\geq 0}$ be the number of logical links from PoD $i$ to $j$ and each PoD $i\in V$ has $R_{i}$ ports. Similar to Google [4], we assume traffic is routed via at most two hops. We define a path as a triad $(s,k,d)$, where $s$ is the source, $d$ is the destination, and $k$ is either an intermediate relay (if $k\neq d$) or the destination itself (if directly routed). The set of all permissible paths is denoted $\mathcal{P}$, and for each source-destination pair $(i,j)$, we define the set of valid relay options as $\mathcal{K}_{ij}=\{k\mid(i,k,j)\in\mathcal{P}\}$. We introduce $f_{i,j,k}$ to denote the fraction of $i\to j$ traffic routed through node $k$ (with $k=j$ representing direct routing). Our goal is to jointly optimize the integer-valued logical topology $n_{i,j}$} and the continuous routing variables $f_{i,j,k}$} to minimize the maximum link utilization $u$. The joint topology and routing optimization (TRO) problem is formulated as Equation 1.
Constraint (1b) limits the total number of logical links per PoD by available port count. Constraint (1c) enforces link symmetry, and requires that $n_{i,j}\in\mathbb{Z}_{\geq 0}$, i.e., each logical link count must be a non-negative integer. Constraints (1d)–(1e) restrict routing to valid two-hop paths. Constraint (1f) bounds link load by capacity. While our framework naturally generalizes to longer path-based routing [8], we adopt the two-hop model for clarity and tractability.
@@ II-B Limitations of Relaxation-Based Approaches
Figure 1: Illustration of infeasibility in the rounded TRO solution produced by COUDER. Assume each PoD has $R_{i}=4$ ports. Subfigure (b) shows a relaxed LP solution with fractional link allocations. Subfigure (c) demonstrates an infeasible rounding outcome where PoD A and B are disconnected. Subfigure (d) shows a feasible and optimal integer solution.
@@ II-C Key Insight and Theoretical Discussion
  • Modular decomposition. ATRO decouples topology and routing to simplify the joint optimization, allowing each subproblem to be solved independently via specialized techniques. This modularity naturally generalizes to multi-hop routing: while our TO component optimizes links based on aggregate loads, the RO component can seamlessly integrate any standard multi-hop TE algorithm.
@@
  • Synergy and Robustness. ATRO converges to near-optimal solutions through the synergy of its components: TO reshapes the global topology while RO mitigates local congestion. In rare sparse-demand cases, ATRO may retain links that an exact solver would prune. This occurs because the routing step lacks an incentive to drive specific link loads to absolute zero, preventing the topology step from removing them. While slightly suboptimal in link count, this behavior enhances robustness by maintaining additional path diversity.
@@ IV-B One-Hop Evaluation (TO Component)
Stress Test. ATRO outperforms MILP-TO in one-hop TO problems, highlighting the need for speed in real-time topology reconfiguration despite MILP-TO’s capability to handle moderate sizes quickly. Stress testing with larger topologies (see Table II) shows ATRO maintains optimal MLU (see Fig.6(a)) and near-second runtimes (see Fig.6(b)), unlike MILP-TO, which struggles with large-scale topology. These findings demonstrate ATRO’s scalability and suitability for latency-sensitive scenarios in one hop scenario.
@@
Link Count Analysis. In addition to MLU and runtime, we evaluate the number of logical links each method provisions to satisfy traffic demand. As seen in Fig. 7, ATRO consistently provisions the fewest links, thanks to theorem 1. It minimizes the consumption of limited OCS port resources, leaving more ”free” ports to accommodate sudden traffic bursts or background flows. MILP-TO often over-provisions due to solver rounding effects, leading to inefficiencies. MCF and BvN lack explicit link minimization objectives and hence result in inflated configurations, particularly in larger topologies like Topo 64 and 128.
@@ VI Conclusion
We present ATRO, a modular framework for computing logical topologies in reconfigurable data center networks (DCNs). In the general multi-hop setting, ATRO alternates between topology optimization (TO) and routing optimization (RO) using lightweight, scalable subroutines. The TO step is solver-free and solved optimally via our proposed Accelerated Binary Search Method (ABSM), while the RO step supports both LP solvers and TE accelerators—enabling fully solver-free execution if desired. Extensive experiments show that ATRO matches or outperforms existing baselines in both performance and runtime, achieving low-latency scheduling in one-hop settings and efficient scalability in large-scale, multi-hop scenarios. Its convergence-guaranteed, plug-and-play design with warm-start and hybrid support makes ATRO well-suited for real-time, dynamic DCNs. In future work, we will enhance ATRO’s robustness to traffic variations.